# Correcting systematic errors in the likelihood optimization of underdamped Langevin models of molecular dynamics trajectories

### v1

@@ II Theory/Methods
Table 1: Mean first passage time (MFPT) for the different Langevin model, inferred using the likelihood correction, computed using Euler-Maruyama integrator, compared to the brute-force MFPT from all-atom MD simulations. Statistical uncertainties are estimated by block averaging.
Taken all together, these results demonstrate that the proposed correction to likelihood inference eliminates systematic errors originating from finite-difference CV velocity estimation: the new scheme is able to parametrize accurate underdamped Langevin models of real complex systems projected on a CV, even for barrier-crossing processes, despite of the use of short, non-ergodic trajectories. We underline the importance of the diagnostic tests we put forward to assess the pertinence and expected accuracy of the model in realistic scenarios, where the brute-force free-energy landscape or kinetic rate are not known beforehand.
The main limitation to the applicability of this method, being based on the Markovian Langevin equation, arises from the existence of long-time memory kernels in several realistic systems. To go beyond this hurdle, two possible strategies could be explored: i) increasing the dimensionality of the model to mitigate non-Markovian effects, or ii) adapting the approach developed in this work to the inference of generalized Langevin models, where memory effects are reproduced by the addition of hidden variables, based on expectation-maximization likelihood schemes [34].
III Acknowledgements
This work was performed with the support of the Institut des Sciences du Calcul et des Données (ISCD) of Sorbonne University (IDEX SUPER 11-IDEX-0004). Calculations were performed on the GENCI-IDRIS French national supercomputing facility, under grant numbers A0110811069, A0130811069. We gratefully acknowledge Line Mouaffac and Jeremy Diharce for suggestions related to code development and simulations, as well as Pierre Ronceray for insightful discussions.

### v2

@@ Abstract
Table 1: Mean first passage time (MFPT) for fullerene dissociation predicted by the Langevin models, inferred using the likelihood correction, computed using Euler-Maruyama integrator, compared to the brute-force MFPT from all-atom MD simulations. Statistical uncertainties are standard deviations of the mean.
In summary, the results presented in this paper demonstrate that the proposed correction to likelihood inference eliminates systematic errors originating from finite-difference CV velocity estimation. The new scheme is able to parametrize accurate underdamped Langevin models of real complex systems projected on a CV, even for barrier-crossing processes, despite of the use of short, non-ergodic trajectories. We underline the importance of the diagnostic tests we put forward to assess the pertinence and expected accuracy of the model in realistic scenarios, where the brute-force free-energy landscape or kinetic rate are not known beforehand. The main limitation to the applicability of this method, based on the Markovian Langevin equation, arises from the existence of long-time memory kernels in several realistic systems. To go beyond this hurdle, two possible strategies could be explored: i) increasing the dimensionality of the model to mitigate non-Markovian effects, or ii) adapting our approach to the inference of generalized Langevin models, where memory effects are reproduced by the addition of hidden variables, based on expectation-maximization likelihood schemes [34].
I Acknowledgements
This work was performed with the support of the Institut des Sciences du Calcul et des Données (ISCD) of Sorbonne University (IDEX SUPER 11-IDEX-0004). Calculations were performed on the GENCI-IDRIS French national supercomputing facility, under grant numbers A0110811069, A0130811069. We gratefully acknowledge Line Mouaffac and Jeremy Diharce for suggestions related to code development and simulations.